# Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach

### v1

@@ 5.2 Structural change and employment growth, wage inequality, and functional distribution of income
In this section we report the results of the estimation of equation 8 using the different outcome variables for employment growth, wage inequality and functional distribution of income. It should borne in mind that, at the present stage, the results presented below aim exclusively at describe the correlation between structural change and the outcome variables, without any implication for their causal relationships.
@@
Table 1 summarises the relationship between structural change ($between$) and employment growth. Looking at columns (1) and (4), the between component of the LWF structural decomposition shows a negative correlation with employment growth at the country level. This indicates that when countries experience a labour flow from low- to high-complexity industries, they also experience slower rates of employment growth. This finding seems to corroborate the hypothesis that high-complexity industries contribute less to aggregate employment, due to their higher levels of knowledge intensity, and lower labour-intensity, as they require fewer (and likely more skilled) workers. It is worth mentioning that neither the within component of the variation in LWF (column 2), nor the sum of both components (column 3) do not show any significant correlation with employment growth, highlighting the importance of structural change in explaining outcomes related to the creation of employment. The coefficient on the between component remains significant also when the within component is also included in the regression.
We now move to the relationship between structural change and wage inequality. First, we look at the link between structural change and the wage gap between top (9th decile) and bottom (1st decile) wages. The wage gaps appears to contract as a result of labour reallocation towards knowledge-intensive industries, as indicated by columns (1) and (4) in Table 2. However, it is not clear whether this result is driven by lower wages in the top of the distribution, or higher wages at the bottom. In order to further explore this, Tables 3 and 4 show respectively the results for the wage gap between 5th/1st and 9th/5th deciles. Interestingly, the wage gap between median and bottom salaries narrows (columns 1 and 4 in Table 3) as a result of structural change, while the ratio between salaries at the 9th and 1st decile doesn’t show any association with structural change (Table 4, columns 1 and 4). We conclude that the reduction in the wage gap between top and bottom wages results from bottom wages growing at a faster rate than top wages. In order to rationalise this result, it is worth going back to the specialisation matrix described in figure 1. As countries move to more complex industries, they drop the least complex ones, as indicated by the "emptier" top-left corner of the matrix. Assuming a positive correlation between complexity and average wages, it can be inferred that the reduction in wage inequality is driven by the fact htat the low-paid jobs in the least complex sectors disappear, making the 1st decile of the salary distribution go up. Although this conjecture is yet to be tested at the present stage of the research, it represents one of the next steps in the analysis conducted in this paper.
Finally, Table 5 displays the results of regressing labour share (the ratio between aggregate wages and value added in the economy) on the between component of LWF. Columns (1) and (4) show a positive relationship between the two, indicating that as workers move to more complex industries, that contributes positively to the increase of the labour component of production, which appears to become more intensive. In order to understand this results, we should rely once more on the assumption that high-complexity industries pay, on average, higher wages. Bearing in mind the diversification patterns displayed by figure 1, coupled with the results on employment growth in Table 1. If employment growth slows down as a result of structural change towards knowledge-intensive industries, the increase in the labour share of the economy is due to be driven by higher average wages rather than higher participation of the population to the workforce. This hypothesis will be tested in the future steps that we will take to carry on with the analysis.
As a final note on the preliminary results presented in this section, it is important to stress that the structural change component is the only one to be significant across the regression specifications summarised by equation 8. As a matter of fact, neither the within component nor the change in labour-weighted fitness explain any of the outcomes observed, indicating that framing structural change as the movement of labour across industries of different complexity – that is, knowledge intensity – can contribute significantly to explain the current trends in employment growth, wage inequality and functional distribution of income in European countries.
@@
[…]
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@
Dependent Variable: Labour share at constant 2015 USD
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@ 5.3 Robustness checks
Given the small number of countries in the sample, we take preliminary steps to confirm that the results illustrated in the previous subsection are not driven by individual countries exhibiting extreme structural change and/or outcome values. In order to achieve this, we perform a set of leave-one-out regressions; for each outcome variable (employment growth, wage inequality and labour share) we run $N=30$ regressions, excluding in each one a different country. The results of this exercise are provided by figure 5, which shows the coefficient on the between component for each of these regressions. As shown in the figure, the coefficient remains always significant, indicating that excluding any country from the sample does not affect the results of any of the models estimated following equation 8.
@@
Additionally, we are concerned that the signal captured by the structural change component (which is a composite measure combining changes in labour share within industries and their complexity) in explaining the outcomes of interest is driven by either of the elements that feed into the index. In order to test whether it is the case or not, we create a different structural change index, which uses the Shannon Entropy ($S$) instead of Complexity:
$LWS_{t}=\sum_{i=1}^{n}\theta_{i,t}S_{i,t};\qquad S=H(X)=-\sum_{i=1}^{n}p(x_{i})\,log\,p(x_{i})$ (12)
@@
[…]
@@
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
We proceed by regressing the newly created index – which is the between component of a different measure although related to LWF, named here Labour-Weighted Entropy – on employment growth, wage inequality, and labour share. Table 6 displays the results of this second robustness exercise, and indicates that the new measure of structural change does a very poor job in explaining the outcomes of interest. This results provides a further confirmation that the complexity-based structural change measure proposed in this paper is able to capture a strong signal that explains outcomes in employment creation and inequality, and proves to be a solid way to conceptualise structural change as the reallocation of workers across industries with different degrees of knowledge intensity.
@@ 6 Conclusions
The empirical analysis conducted in this paper has offered insights on the relationship between structural change – framed as the reallocation of labour towards more complex and knowledge intensive activities – and employment growth, wage inequality, and functional distribution of income. This relationship has been examined introducing several innovative methodological and analytical elements. In the first place, we have introduced a novel method to infer comparative advantages in an unbiased way, in what we believe to be the first empirical application of the BiWCM method proposed by Bruno et al. (2023). Secondly, we have constructed measures of Economic Fitness and Complexity that allow the comparison of these two indices over time (Mazzilli et al. 2024), overcoming a known limitation in the use of EC indices in econometric models (Caldarola et al. 2023). Third, we have introduced a new measure – the Labour-Weighted Fitness – whose variation allows to be decomposed in such a way as to capture the reallocation of labour across industries with different levels of complexity. This strategy has proved to have significant explanatory power with respect to the creation of employment, the distribution of wages, and the share of labour participation to the aggregate value added.
@@
The preliminary results indicate that structural change towards complex industries slows down employment creation, due to the lower labour intensity of complex industries. However, the gap between top and bottom wages has shrunk, offering new insights on the inequality reduction brought about by structural change. It must be noted that wage inequality reduces as a result of an increase of the average wages at the first decile of the wage distribution, which in turn is the result of the loss of low-complexity (and likely low pay) jobs by the fittest most diversified countries. The increase in average wage also helps explain the relationship between structural change and the labour share, with the latter increasing when workers move to high-complexity industries. Once more, this effect is not likely to by ascribable to a higher rate of participation to the active workforce, but by the higher salaries paid to workers in countries specialised in complex industries. As a further step to confirm our results, we plan to examine the relationship between the structural change component of the LWF fitness and average wages, which may come in help to rationalise our findings on wage and functional distribution of income. Moreover, in order to confirm the robustness of the results, we are planning to devise an identification strategy based on a shift-share instrument, which appears at the moment the most suitable identification strategy given the nature of our (potentially) endogenous explanatory variable of interest.
Our results bear important consequences for policy making, as they highlight a trade-off between industrial upgrading and labour creation, while also dissecting the nature of the reduction of wage and functional inequality resulting from the movement of labour towards complex industries. Based on these results, policy-makers will have to device policy tools that steer structural change towards those industries that, at the same time, conflate job creation with higher technological sophistication, in order to sustain a process of inclusive economic growth. Such industries can be identified using the EFC framework with a focus on labour, rather than exports alone. Coupling the geography of productive capabilities with labour and inequality, enables us to advance with respect to the extant literature not only in providing a policy-relevant and data-driven toolkit to embrace socio-economic complexity, but also in offering a novel analytical and empirical understanding of a pathway not only geared towards boosting growth, but also to steering inclusive and sustainable development.

### v2

@@ 5.2 Structural change and employment growth, wage inequality, and functional distribution of income
In this section we report the results of the estimation of Equation 7 using the different outcome variables for employment growth, wage inequality and functional distribution of income. It should borne in mind that the results presented below aim exclusively at describing the correlation between structural change and the outcome variables, without any implication for their causal relationships.
@@
Table 1 summarises the relationship between structural change (the variable Between in the table) and employment growth. Looking at columns (1) and (4), the between-component of the LWF structural decomposition shows a negative correlation with employment growth at the country level. This indicates that when countries experience a labour flow from low- to high-complexity industries, they also experience slower rates of employment growth. This finding seems to corroborate the hypothesis that high-complexity industries contribute less to aggregate employment growth, due to their higher levels of knowledge intensity, and lower labour intensity, as they require fewer (and likely more skilled) workers. It is worth mentioning that neither the within-component of the variation in LWF (column 2), nor the sum of both components (column 3) shows any significant correlation with employment growth, highlighting the importance of structural change in explaining outcomes related to the creation of employment. The coefficient on the between-component remains significant also when the within-component is also included in the regression.
@@
Dependent Variable: Yearly employment rate growth ($\%$)
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variable is the growth rate in employment between year $t$ and t1t-1 . Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Fitness decomposition (both from Equation 6). $\Delta LWF$ denotes the change in the Labour Weighted Fitness index. GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
We now move to the relationship between structural change and wage inequality. First, we look at the link between structural change and the top-to-bottom inequality, i.e. the wage gap between top (9th decile) and bottom (1st decile) wages. As can be appreciated by the negative and significant coefficients of the between-component (alone and together with the within-component respectively in column 1 and column 4) in Table 2, the 9th/1st wage ratio appears to contract as a result of labour reallocation towards knowledge-intensive industries. However, it is not clear whether this result is driven by lower wages in the top of the distribution, or higher wages at the bottom. To further explore this, Tables 3 and 4 show the results for upper- and lower-tail inequality, the 5th/1st and 9th/5th wage ratios, respectively. Interestingly, structural change is associated with a narrowing of the wage gap between median and bottom wages (columns 1 and 4 in Table 3), whereas it does not appear to be significantly associated with the wage gap between top and median wages (columns 1 and 4 in Table 4).
The combination of these results, therefore, suggests that the reduction in the wage gap between top and bottom wages is driven mainly by changes at the bottom of the wage distribution. In order to rationalise this result, it is worth going back to the specialisation matrix described in Figure 1. As countries move to more complex industries, they drop the least complex ones, as indicated by the "emptier" top-left corner of the matrix. Assuming a positive correlation between complexity and average wages, it can be inferred that the reduction in wage inequality is driven by the fact that the very low-paid jobs in the least complex sectors disappear, making the 1st decile of the salary distribution go up.
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variable is the wage ratio between the 9th and 1st deciles of the wage distribution. Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Fitness decomposition (both from Equation 6). $\Delta LWF$ denotes the change in the Labour Weighted Fitness index. GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variable is the wage ratio between the 5th and 1st deciles of the wage distribution. Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Fitness decomposition (both from Equation 6). $\Delta LWF$ denotes the change in the Labour Weighted Fitness index. GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variable is the wage ratio between the 9th and 5th deciles of the wage distribution. Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Fitness decomposition (both from Equation 6). $\Delta LWF$ denotes the change in the Labour Weighted Fitness index. GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Finally, Table 5 displays the results of regressing the labour share (defined as the ratio between aggregate wages and value added in the economy) on the between-component of LWF. Columns (1) and (4) show a positive relationship between the two, indicating that, as workers move to more complex industries, the labour share tends to increase. This indicates that structural change towards more complex industries is associated with a higher share of labour in total value added. To interpret this finding, we should rely once more on the evidence that high-complexity industries pay, on average, higher wages, as shown by Sbardella et al. (2017); Aufiero et al. (2024). Bearing in mind the diversification patterns displayed by Figure 1, coupled with the results on employment growth in Table 1, if employment growth slows down as a result of structural change towards knowledge-intensive industries, the increase in the labour share of the economy is due to be driven by higher average wages rather than broader employment expansion. This positive labour share-structural change association is particularly noteworthy because it contrasts with the broad declining trend in the labour share documented for European economies over recent decades (Arpaia et al., 2009; Stockhammer, 2012; Dimova, 2019). While at the aggregate level the literature has generally associated this decline with capital deepening, technological change, globalisation, and weakening labour market institutions, our findings suggest that, holding other factors constant, structural upgrading towards more complex sectors appears as a countervailing force to the aggregate downward trend, suggesting that this specific component of structural transformation may partly offset the broader pressures that have contributed to the decline in the labour share. This positive effect is also consistent with what observed by Hubmer and Restrepo (2026), who document a rise in the labour share of typical firms, showing how larger firms are the ones that drive the capital-labour substitution driven by automation. An expansion of industries dominated by firms similar to the typical firm would be associated also to a growth in the labour share.
@@
Dependent Variable: Labour share (constant 2015 USD)
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variable is the share of labour compensation over Gross Domestic Product, both measured in constant 2015 US Dollars. Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Fitness decomposition (both from Equation 6). $\Delta LWF$ denotes the change in the Labour Weighted Fitness index. GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Relatedly, and as a final note on the results presented in this section, it is important to stress that the structural change component is the only one to be significant across all the regression specifications summarised by Equation 7. As a matter of fact, neither the within-component nor the change in labour-weighted fitness explain any of the outcomes observed, indicating that framing structural change as the movement of labour across industries of different complexity can contribute significantly to explain the current trends in employment growth, wage inequality and functional distribution of income in European countries.
@@ 5.3 Robustness checks
Given the small number of countries in the sample, our econometric exercise is an inherently small N one, therefore sensitive to changes in individual observations. We test whether the results illustrated in the previous subsection are driven by individual countries exhibiting extreme structural change and/or outcome values. We perform a set of leave-one-out regressions: for each outcome variable (employment growth, wage inequality, and labour share) we run $N=30$ regressions, excluding in each one a different country. The results of this exercise are provided by Figure 5, which shows the coefficient of the between-component for each of these regressions. As shown in the figure, the coefficient remains always significant, indicating that excluding any country from the sample does not affect the results of any of the models estimated following Equation 7.
@@
  • Notes: The figure reports the coefficients on the structural change component (the between-term of the Labour-Weighted Fitness decomposition, as defined in Equation 6) obtained from leave-one-out regressions for each of the three outcome variables: yearly employment rate growth (top-left panel), the 9th-to-1st decile wage ratio (top-right panel), and the labour share of the economy (bottom-left panel). Each panel reports the results of N = 30 separate regressions, in each of which one country is excluded from the sample. All specifications follow Equation 7 and include country and year fixed effects, the within-component of the LWF decomposition, log GDP per capita, log population, R&D expenditure as a share of GDP, and exports as a share of GDP. Standard errors are clustered at the country-year level. Vertical bars represent 90% confidence intervals around each point estimate. The horizontal black line denotes the zero threshold. Each coloured line corresponds to a regression in which a different country has been excluded from the sample.
Additionally, a potential concern is that the signal captured by our structural change component (which is a composite measure combining changes in labour share within industries and their complexity) in explaining the outcomes of interest is driven by either of the elements that feed into the index. In order to test whether it is the case or not, we construct an alternative structural change index, in which the Complexity of industries is replaced by Shannon Entropy. For each time $t$, we compute a Labour Weighted Shannon entropy indicator ($LWS_{t}$) as follows:
$LWS_{t}=\sum_{i=1}^{n}\theta_{i,t}S_{i,t};\qquad S=H(X)=-\sum_{i=1}^{n}p(x_{i})\,log\,p(x_{i}).$ (11)
LWS is a measure of sectoral dispersion, as it captures how employment is distributed across industries and therefore the degree of sectoral diversification of the economy. However, it is agnostic about the qualitative content of such diversification: a increase in entropy may reflect a more even distribution of labour across industries, but does not distinguish whether workers are moving into more or less complex industries. In contrast, the Economic Fitness and Complexity framework is explicitly designed to capture the asymmetric distribution of productive capabilities across industries and countries. Accordingly, LWF does not simply reflect whether labour is becoming more dispersed across activities, but rather captures the direction of structural change, identifying whether workers are being reallocated in more complex, capability-intensive sectors. In other words, while entropy captures diversification and thus the breadth of the productive structure, LWF reflects the direction of structural change by identifying movement towards more complex industries and therefore its position in the hierarchy of productive sophistication.
@@
Dependent Variables: Emp. growth ($\%$) Ratio $9^{th}/1^{st}$ dec. Labour share
@@
  • Notes: Clustered (Country-Year) standard errors reported in parentheses. All specifications include country and year fixed effects. The dependent variables are: the growth rate in employment between year $t$ and t1t-1 (1), the wage ratio between the 9th and 1st deciles of the wage distribution (2), and the share of labour compensation over Gross Domestic Product, both measured in constant 2015 US Dollars (3). Between captures the between-component of the yearly change in Labour Weighted Fitness, while Within captures the within-component of the Labour Weighted Entropy decomposition (both from adapting Equation 6). GDPpc (log) is log GDP per capita in US dollars (constant 2015); Population (log) is log total population; R&D ($\%$ GDP) is gross domestic expenditure on research and development as a percentage of GDP; Exports ($\%$ GDP) is total exports as a percentage of GDP. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
We proceed by regressing the between-component of $LWS$ on employment growth, wage inequality as proxied by the different wage ratios described above, and the labour share. Table 6 displays the results of this second robustness exercise, and indicates that the new entropy-based measure of structural change does a very poor job in explaining all three outcomes of interest. This result is informative because it shows how the two measures (fitness and entropy) capture different aspects of structural change. In addition, it corroborates that the explanatory power of the complexity-based structural change measure proposed in this paper does not stem from a generic sectoral compositional change, but from its ability to capture structural upgrading, and is thus is able to capture a strong signal that explains outcomes in employment creation and inequality. In sum, our measure proves to be a solid way to conceptualise structural change as the reallocation of workers across industries with different degrees of complexity.
@@ 6 Conclusions
The empirical analysis conducted in this paper has offered insights on the relationship between structural change – framed as the reallocation of labour towards more complex industries – and employment growth, wage inequality, and functional distribution of income. The results indicate that structural change towards complex industries slows down employment creation, due to the lower labour intensity of complex industries. However, the gap between top and bottom wages has shrunk, offering new insights on the inequality reduction brought about by structural change. It must be noted that wage inequality reduces as a result of an increase of the average wages at the first decile of the wage distribution, which in turn is the result of the loss of low-complexity (and likely low pay) jobs by the fittest most diversified countries. The increase in average wage also helps explain the relationship between structural change and the labour share, with the latter increasing when workers move to high-complexity industries. Once more, this effect is not likely to by ascribable to a higher rate of participation to the active workforce, but by the higher salaries paid to workers in countries specialised in complex industries. As a further step to confirm our results, we plan to examine the relationship between the structural change component of the LWF fitness and average wages, which may come in help to rationalise our findings on wage and functional distribution of income. Moreover, in order to confirm the robustness of the results, we are planning to devise an identification strategy based on a shift-share instrument, which appears at the moment the most suitable identification strategy given the nature of our (potentially) endogenous explanatory variable of interest.
This relationship has been examined introducing several innovative methodological and analytical elements. In the first place, we have introduced a novel method to infer comparative advantages in an unbiased way, in what we believe to be the first empirical application of the BiWCM method proposed by Bruno et al. (2023). Secondly, we have constructed measures of Economic Fitness and Complexity that allow the comparison of these two indices over time (Mazzilli et al., 2024), overcoming a known limitation in the use of EC indices in econometric models (Caldarola et al., 2024). Third, we have introduced a new measure – the Labour-Weighted Fitness – whose variation allows to be decomposed in such a way as to capture the reallocation of labour across industries with different levels of complexity. This strategy has proved to have significant explanatory power with respect to the creation of employment, the distribution of wages, and the share of labour participation to the aggregate value added.
Our findings carry several implications for the design of innovation and industrial policy within the European context. First, they suggest that policies aimed exclusively at promoting structural upgrading—through R&D subsidies, technology adoption incentives, or sector-specific support for knowledge-intensive activities – may deliver gains in productivity and average wages, but at the cost of slower employment creation. This trade-off cannot be resolved by innovation policy alone: it requires explicit coordination between innovation, industrial, and labour-market policies, in line with recent calls for a more integrated and mission-oriented approach to economic transformation. Second, the heterogeneity in labour absorption across industries that emerges from our analysis indicates that not all forms of structural upgrading are equivalent from a distributive standpoint. The EFC framework, when applied to employment rather than trade data alone, can serve as a diagnostic tool to identify activities that combine high complexity with substantive labour absorption. Third, our finding that wage compression at the bottom of the distribution is largely driven by the disappearance of low-complexity jobs implies that the inclusiveness of structural change is contingent on what happens to displaced workers: in the absence of effective active labour-market policies, reskilling programmes, and social protection, the same process that compresses measured wage inequality may simultaneously generate non-employment and labour-market exit among low-skilled workers, with the observed reduction in inequality reflecting selection out of employment rather than genuine upward mobility. Our findings thus argue for a policy mix that moves beyond stimulating capability accumulation per se, and instead deliberately steers structural transformation towards technological trajectories that conflate sophistication with employment-generating potential, while building the institutional infrastructure – labour standards, collective bargaining, training systems, and welfare provisions – needed to ensure that the gains from upgrading are broadly shared.